# FoC: Figure out the Cryptographic Functions in Stripped Binaries with LLMs

### v2

@@ 4. Methodology Overview
In this section, we provide an overview of the FoCframework, as illustrated in Figure 4. The framework comprises two key components: (1) Binary Large Language Model (FoC-BinLLM), which is designed for understanding the pseudo-code of cryptographic functions that are lifted from stripped binaries. FoC-BinLLM predicts the stripped function names and generates concise summaries of the functions. (2) Binary Code Similarity Model (FoC-Sim), which leverages the semantic encoder from FoC-BinLLM and also incorporates graph structure information with Graph Convolutional Network (GCN), along with cryptographic features. These components are briefly described below.
@@
We adopt the Transformer model following the encoder-decoder architecture, enabling flexible use in either encoder-only mode for semantic embedding generation, or in encoder-decoder mode for causal generation. As illustrated in Figure 4 ①, we initialize our model with the pre-trained weights from a gold implementation, allowing us to avoid heavy training from scratch. The base model is then fine-tuned on our cryptographic binary dataset, optimizing it specifically for binary code understanding. We train the base model on our cryptographic binary dataset to specialize it for understanding binary code. Our FoC-BinLLM processes the pseudo-code of an unknown function as input and generates a formatted summary for analysts, providing detailed semantic insights and mitigates challenge C3. Additionally, we propose a keyword-based discriminator to determine the class of critical behavior of cryptographic functions based on model predictions.
@@
As shown in Figure 4 ②, FoC-Sim utilizes the pseudo-code, the attributed control-flow graph (ACFG), and the cryptographic features as inputs, which can be easily extracted from the binary function using a modern decompiler, such as IDA Pro (IDA_PRO, ). The model then creates an embedding representation of the function. It is aware of any changes in the binary code, which compensates for the lack of sensitivity of our generative model and mitigates the challenge C4. By leveraging FoC-Sim, we can access source-level information when analyzing cryptographic algorithms in binaries, provided that homologous implementations are present in the database. Additionally, FoC-Sim can be employed to inspect binaries for identifying vulnerable cryptographic implementations.
@@ 5.1. Binary Large Language Model (FoC-BinLLM)
CodeT5+ is initialized with weights from previous pre-trained LLMs (i.e., CodeGen-mono (CodeGen_ICLR_2023, )), and is trained on two large-scale source code datasets: a multilingual dataset  [^2] https://huggingface.co/datasets/codeparrot/github-code containing 51.5 billion tokens, and the CodeSearchNet released by previous research (CodeSearchNet_CoRR_2019, ). Benefiting from the training on bimodal data consisting of both natural language (NL) and programming language (PL), CodeT5+ achieved state-of-the-art performance in various downstream tasks, such as code generation and code summarization, at the time of its release. The pre-built understanding and generation capabilities of CodeT5+ in both NL and PL form the foundation for training it to become an expert LLM.
@@ 7.1. Answer to RQ1: Performance in Cryptographic Binary Code Summarization
Results Analysis. Table 8 shows the detailed performance. FoC-BinLLM performs impressively on the test set. Specifically, it achieves scores of 41.34%, 14.47%, and 40.20% on the ROUGE-L, BLEU-4, and METEOR metrics, respectively, outperforming all existing baseline methods. This demonstrates that FoC-BinLLM has higher accuracy and expressiveness in generating natural language summaries of cryptographic binary code, and can effectively capture the key semantics of binary functions.
The expert models BinT5 and HexT5, both exhibit a significant performance degradation compared to the results reported in their articles, which may be attributed to our cryptographic binary surpassing the domain where they collect their dataset, especially in the complexity of the cryptographic algorithm. In contrast, FoC-BinLLM is able to handle these challenges better, outperforming 29.72%, 13.32%, and 30.29% on average in the three metrics, showing its unique advantages in the cryptographic domain task. Compared with general LLMs such as Mixtral and ChatGPT, FoC-BinLLM also performs well in three key metrics, outperforming them by 3.28%, 2.15%, and 5.59% on average respectively. It is worth noting that FoC-BinLLM has only 220M parameters, while Mixtral (8x7B) and ChatGPT have much larger scales than it, but show significant limitations in cryptographic binary codes. This comparison reflects that customized training of models for specific domain tasks can often effectively improve model performance.
@@ 7.2.1. Performance on the General Dataset
7.2.2. Performance on the Cryptographic Dataset
Baselines. To further evaluate, in addition to the baselines used in the benchmark in Section 7.2.1, we also employ recent state-of-the-art BCSD methods, i.e., PalmTree (BCSD_PalmTree_CCS_2021, ), Trex (BCSD_TREX_TSE_2023, ), jTrans (BCSD_jTrans_ISSTA2022, ). PalmTree is based on BERT and performs self-supervised pre-training on large-scale unlabeled binary corpora to generate instruction embeddings that can be used to calculate function similarity. Trex utilizes transfer-learning-based models to learn execution semantics from micro-traces to generate function embeddings. jTrans manages to embed control-flow information into Transformer and pre-trains it on a large-scale dataset. Especially, we did not evaluate the performance of PalmTree on XA sub-task, because it only supports the x86 architecture.
Results Analysis. The results are shown in Table 11. FoC-Sim also demonstrates excellent performance on cryptographic dataset. Specifically, in the One-to-one Comparison scenario, FoC-Sim achieves over 99% AUC on all four sub-tasks, surpassing the best-performing baseline, jTrans, by 6.7%, 7.3%, 15.3%, and 15.3%, respectively. Especially in the One-to-many Search scenario, FoC-Sim excelles with 94.0% MRR@10, 91.0% Recall@1, and 99.0% Recall@10 on the XM sub-task, significantly higher than the baseline jTrans, which achieves only 46.3%, 38.0%, and 66.8%.
FoC-Sim also offers high inference efficiency. As indicated in Table 11, FoC-Sim requires an average of 62.26ms to perform 100 function similarity comparisons, which includes the model inference and similarity score calculation process. This time is comparable to methods like SAFE and PalmTree, and significantly outperforms GMN, Asm2Vec, jTrans, and Trex. Although Zeek achieves a faster time of 28.14ms, considering the substantial performance improvements offered by FoC-Sim, the slight increase in time cost is a reasonable trade-off.
@@ 7.3. Answer to RQ3: Performance in Practical Ability
In this section, we explore the performance of FoC  in two real-world scenarios: (1) analyzing cryptographic functions in a virus, and (2) retrieving vulnerable cryptographic implementations in firmware.
@@ 7.3.1. Cryptographic Function Analysis in Virus
7.3.2. Vulnerablity Detection in Firmware
@@
To further explore the practical capabilities of our similarity model FoC-Sim, we utilize it to detect vulnerable cryptographic functions in binaries. To do this, we obtain the firmware to create a firmware database and compile the vulnerable functions and their patched functions to create a vulnerability database. Subsequently, we perform two search tasks. (1) Utilizing vulnerability database, we apply FoC-Sim to detect vulnerable cryptographic functions in the firmware database. It is considered a successful identification if the vulnerable functions are found among the top 10 most similar in all functions from a suspicious file. (2) We evaluate the ability of FoC-Sim to distinguish between vulnerable functions and patched functions. Specifically, a vulnerable function from the firmware database is considered to be successfully distinguished if it has a higher similarity to the vulnerable version rather than the patched version.
@@
Results Analysis. Table 12 shows the detailed experimental results. The ”Vulnerability Detection” column in Table 12 demonstrates that FoC-Sim is capable of accurately detecting the majority of vulnerable functions across the firmware of different vendors. Specifically, it identified 91 out of 97 functions in NETGEAR firmware, 168 out of 181 in TP-LINK firmware, and 319 out of 330 in Xiaomi firmware. As shown in the “Vulnerability/Patch Distinction” column, FoC-Sim can correctly distinguish vulnerable functions from patched functions. It correctly distinguishes 70 out of 97 functions in NETGEAR firmware, 142 out of 181 in TP-Link, and 289 out of 330 in Xiaomi. It demonstrates that FoC-Sim has the ability to overcome the challenge C4. However, we observe a failed case in Libgcrypt. With manual inspection, we find that there is a huge difference between the vulnerable functions from the firmware database and the vulnerability database in both text and structure. In other words, the vulnerable function in Libgcrypt has multiple significantly different versions, while our vulnerable database only includes one.
@@
Answering RQ3: In the process of analyzing the encrypted communication module in the virus sample, FoC-BinLLM accurately predict the function names and summaries of related functions. FoC-Sim can effectively retrieve potential vulnerable cryptographic functions in the firmware, and correctly distinguish the subtle differences between the vulnerable functions and the patched versions. In summary, the experimental results show the practical ability of FoC  in real-world scenarios.

### v3

@@ 4. Methodology Overview
In this section, we provide an overview of the FoC framework, as illustrated in Figure 4. The framework comprises two key components: (1) Binary Large Language Model (FoC-BinLLM), which is designed for understanding the pseudo-code of cryptographic functions that are lifted from stripped binaries. FoC-BinLLM predicts the stripped function names and generates concise summaries of the functions. (2) Binary Code Similarity Model (FoC-Sim), which leverages the semantic encoder from FoC-BinLLM and also incorporates graph structure information with Graph Convolutional Network (GCN), along with cryptographic features.
These two components, though performing different tasks, are tightly integrated in the FoC framework to form an efficient collaborative working mechanism. During the training process, FoC-BinLLM is first trained to perform seq-to-seq tasks using an encoder-decoder mode. Subsequently, FoC-Sim directly uses the encoder part trained by FoC-BinLLM to train the embedding generation task. Since the encoder has been fully trained and adapted to the domain knowledge of binary code and cryptographic functions, we freeze its parameters and only train the lightweight GCN and MLP, which greatly improves training efficiency and reduces computational overhead. During inference, FoC-BinLLM and FoC-Sim share the same encoder part, and after leaving the encoder, they perform parallel inference. For an input binary function, the FoC framework only needs to process the input once to simultaneously generate a semantic summary of the function (completed by FoC-BinLLM) and retrieve homologous similar functions (completed by FoC-Sim), and the information they output is complementary to each other for human reverse engineers. The close synergy between the two models provides a more comprehensive understanding of binary cryptographic operations. These components are briefly described below.
@@
We adopt the Transformer model following the encoder-decoder architecture, enabling flexible use in either encoder-only mode for semantic embedding generation, or in encoder-decoder mode for causal generation. As illustrated in Figure 4 ①, we initialize our model with the pre-trained weights from a gold implementation, allowing us to avoid heavy training from scratch. The base model is then fine-tuned on our cryptographic binary dataset, optimizing it specifically for binary code understanding. We train the base model on our cryptographic binary dataset to specialize it for understanding binary code. Our FoC-BinLLM processes the pseudo-code of an unknown function as input and generates a formatted summary for analysts, providing detailed semantic insights and mitigates challenge C2. Additionally, we propose a keyword-based discriminator to determine the class of critical behavior of cryptographic functions based on model predictions.
@@
As shown in Figure 4 ②, FoC-Sim utilizes the pseudo-code, the attributed control-flow graph (ACFG), and the cryptographic features as inputs, which can be easily extracted from the binary function using a modern decompiler, such as IDA Pro (IDA_PRO, ). The model then creates an embedding representation of the function. It is aware of any changes in the binary code, which compensates for the lack of sensitivity of our generative model and mitigates the challenge C3. By leveraging FoC-Sim, we can access source-level information when analyzing cryptographic algorithms in binaries, provided that homologous implementations are present in the database. Additionally, FoC-Sim can be employed to inspect binaries for identifying vulnerable cryptographic implementations.
@@ 5.1. Binary Large Language Model (FoC-BinLLM)
CodeT5+ is initialized with weights from previous pre-trained LLMs (i.e., CodeGen-mono (CodeGen_ICLR_2023, )), and is trained on two large-scale source code datasets: a multilingual dataset  [^2] https://huggingface.co/datasets/codeparrot/github-code containing 51.5 billion tokens, and the CodeSearchNet released by previous research (CodeSearchNet_CoRR_2019, ). Benefiting from the training on bimodal data consisting of both natural language (NL) and programming language (PL), CodeT5+ achieved state-of-the-art performance in various downstream tasks, such as code generation and code summarization, at the time of its release.
From an architectural perspective, the encoder-decoder architecture of CodeT5+ provides it with extremely high flexibility in both understanding and generating tasks. It can flexibly switch between encoder-only, decoder-only, and encoder-decoder modes, perfectly adapting to our diverse needs in binary code analysis. The architectural flexibility of CodeT5+, as well as the pre-built understanding and generation capabilities in both NL and PL, form the foundation for training it to become an expert LLM for binary code tasks.
@@ 7.1. Answer to RQ1: Performance in Cryptographic Binary Code Summarization
Results Analysis. Table 8 shows the detailed performance. FoC-BinLLM performs impressively on the test set. Specifically, it achieves scores of 41.34%, 14.47%, and 40.20% on the ROUGE-L, BLEU-4, and METEOR metrics, respectively, outperforming all existing baseline methods. This demonstrates that FoC-BinLLM has higher accuracy and expressiveness in generating natural language summaries of cryptographic binary code, and can effectively capture the key semantics of binary functions. The expert models BinT5 and HexT5, both exhibit a significant performance degradation compared to the results reported in their articles, which may be attributed to our cryptographic binary surpassing the domain where they collect their dataset, especially in the complexity of the cryptographic algorithm. In contrast, FoC-BinLLM is able to handle these challenges better, outperforming 29.72%, 13.32%, and 30.29% on average in the three metrics, showing its unique advantages in the cryptographic domain task. Compared with general LLMs such as Mixtral and ChatGPT, FoC-BinLLM also performs well in three key metrics, outperforming them by 3.28%, 2.15%, and 5.59% on average respectively. It is worth noting that FoC-BinLLM has only 220M parameters, while Mixtral (8x7B) and ChatGPT have much larger scales than it, but show significant limitations in cryptographic binary codes. This comparison reflects that customized training of models for specific domain tasks can often effectively improve model performance.
@@
Table 9. Comparison with existing methods on function name prediction.
Method # Parameters Metrics Time(s)1
Precision Recall F1-score
SymLM (SYMLM_CCS_2022, ) - 0.103 0.141 0.134 0.183
XFL (patrick2023xfl, ) 377M 0.207 0.152 0.173 0.102
Mixtral (jiang2024mixtral, ) 8x7B 0.179 0.233 0.186 1.801
ChatGPT (chatgpt-GPT3.5-2022, ) - 0.200 0.199 0.195 -
FoC-BinLLM 220M 0.302 0.317 0.326 0.062
  • 1
    Average time cost on each binary function.
Function Name Prediction. Since function name prediction is one of the training tasks of FoC-BinLLM, we evaluate the performance of FoC-BinLLM in comparison with recent related methods, such as SymLM (SYMLM_CCS_2022, ) and XFL (patrick2023xfl, ), on the task. SymLM leverages the pre-trained Trex (BCSD_TREX_TSE_2023, ) model to extract execution flow information and concatenates the function embeddings to capture calling context information. Additionally, it employs the CodeWordNet model to alleviate the problem of ambiguous function names (e.g., synonyms, abbreviations, etc.). XFL introduces and information aggregation strategy by concatenating global and contextual embeddings to preserve both types of information. it also utilizes PfastreXML (jain2016extreme, ) and a binary function embedding to effectively perform multi-label classification of function name tokens. Furthermore, similar to the binary code summarization experiment, we select two large language models, Mixtral-8x7B-Instruct-v0.1 and ChatGPT-3.5-turbo-1106, and apply an in-context learning approach for function name prediction as our baselines.
The results are shown in Table 9. FoC-BinLLM achieves Precision, Recall, and F1-score of 30.2%, 31.7%, and 32.6%, respectively, significantly outperforming baseline methods. This demonstrates its superior capability in binary function name prediction, enabling it to more accurately capture the behavioral characteristics of binary functions. Compared to expert models specifically designed for binary function name prediction such as SymLM and XFL, FoC-BinLLM improves the F1-score by 19.2% and 15.3%, respectively. The relatively poor performance of SymLM and XFL can be attributed to their insufficient generalization capabilities and lack of cryptography domain-specific knowledge. Furthermore, despite having only 220M parameters, which is substantially fewer than the general LLMs Mixtral and ChatGPT, FoC-BinLLM still achieves remarkable performance gains, with an F1-score improvement of 14.0% over Mixtral and 13.1% over ChatGPT. This highlights the value of domain-specific models in binary analysis. Additionally, FoC-BinLLM demonstrates strong efficiency, with an inference time of only 0.062 seconds per function, underscoring its practicality in real-time analysis scenarios.
@@ 7.2.1. Performance on the General Dataset
7.2.2. Performance on the Cryptographic Dataset
Baselines. To further evaluate, in addition to the baselines used in the benchmark in Section 7.2.1, we also employ recent state-of-the-art BCSD methods, i.e., PalmTree (BCSD_PalmTree_CCS_2021, ), Trex (BCSD_TREX_TSE_2023, ), jTrans (BCSD_jTrans_ISSTA2022, ), Asteria-Pro (yang2023asteria, ) and RCFG2Vec (li2024rcfg2vec, ). PalmTree is based on BERT and performs self-supervised pre-training on large-scale unlabeled binary corpora to generate instruction embeddings that can be used to calculate function similarity. Trex utilizes transfer-learning-based models to learn execution semantics from micro-traces to generate function embeddings. jTrans manages to embed control-flow information into Transformer and pre-trains it on a large-scale dataset. Asteria-Pro uses the Tree-LSTM to encode the AST into a representation vector and then reorders the candidate functions using function call relationship. RCFG2Vec uses syntax tree-based instruction embedding technology and acyclic graph neural network to solve the OOV problem and long-distance dependencies between basic blocks in existing methods respectively. Especially, we did not evaluate the performance of PalmTree on XA sub-task, because it only supports the x86 architecture.
Results Analysis. The results are shown in Table 12. FoC-Sim also demonstrates excellent performance on cryptographic dataset. Specifically, in the One-to-one Comparison scenario, FoC-Sim achieves over 99% AUC on all four sub-tasks, surpassing the best-performing baseline, RCFG2Vec, by 2.6%, 2.1%, 6.4%, and 5.2%, respectively. Especially in the One-to-many Search scenario, FoC-Sim excelles with 94.0% MRR@10, 91.0% Recall@1, and 99.0% Recall@10 on the XM sub-task, significantly higher than the baseline RCFG2Vec, which achieves only 77.5%, 69.4%, and 89.1%.
Figure 8. Results of Recall@1 under different pool sizes (XM sub-task in One-to-many Search).
FoC-Sim also offers high inference efficiency. As indicated in Table 12, FoC-Sim requires an average of 62.26ms to perform 100 function similarity comparisons, which includes the model inference and similarity score calculation process. This time is comparable to methods like SAFE and PalmTree, and significantly outperforms GMN, Asm2Vec, jTrans, Trex, Asteria-Pro and RCFG2Vec. Although Zeek achieves a faster time of 28.14ms, considering the substantial performance improvements offered by FoC-Sim, the slight increase in time cost is a reasonable trade-off.
Additionally, as shown in Figure 8, we further analyze the impact of different pool sizes on the Recall@1 in the One-to-many Search scenario. Specifically, we select the five baseline methods that performed well in the XM sub-task in Table 12, namely PalmTree, Trex, jTrans, Asteria-Pro, and RCFG2Vec, and gradually increase the pool size from 10 to 10,000. The experimental results demonstrate that as the pool size increases, the task difficulty rises significantly, leading to a noticeable drop in Recall@1 for all methods. However, FoC-Sim consistently outperforms the baseline methods across all pool sizes. Notably, the performance degradation of FoC-Sim is relatively small as the pool size increases. Even at the maximum pool size of 10,000, FoC-Sim maintains high performance, showcasing its superior adaptability and robustness in handling large-scale and complex real-world scenarios. These results further validate the effectiveness of FoC-Sim.
@@ 7.3. Answer to RQ3: Performance in Practical Ability
In this section, we explore the performance of FoC  in two real-world scenarios: (1) analyzing cryptographic functions in a virus, and (2) retrieving cryptographic implementations in firmware.
@@ 7.3.1. Cryptographic Function Analysis in Virus
7.3.2. Cryptographic Implementation Detection in Firmware
@@
To further explore the practical capabilities of our similarity model FoC-Sim, we utilize it to detect cryptographic implementation functions in binaries. To do this, we obtain the firmware to create a firmware database and compile the vulnerable functions and their patched functions to create a vulnerability database. Subsequently, we perform three search tasks. (1) Utilizing vulnerability database, we apply FoC-Sim to detect vulnerable cryptographic functions in the firmware database. It is considered a successful identification if the vulnerable functions are found among the top 10 most similar in all functions from a suspicious file. (2) We evaluate the ability of FoC-Sim to distinguish between vulnerable functions and patched functions. Specifically, a vulnerable function from the firmware database is considered to be successfully distinguished if it has a higher similarity to the vulnerable version rather than the patched version. (3) We attempt to utilize FoC-Sim to detect cryptographic functions from real-world stripped binaries.
@@
Table 14. Results of detecting cryptographic functions in real-world stripped binaries.
Model Metrics
Recall@1 Recall@10 Recall@100
FoC-Sim 0.714 0.853 0.955
w/o Code Semantics 0.577 0.706 0.843
w/o Control Structure 0.692 0.826 0.923
w/o Cryptographic Features 0.701 0.842 0.931
@@
Results Analysis. Table 13 shows the detailed experimental results. The ”Vulnerability Detection” column in Table 13 demonstrates that FoC-Sim is capable of accurately detecting the majority of vulnerable functions across the firmware of different vendors. Specifically, it identified 91 out of 97 functions in NETGEAR firmware, 168 out of 181 in TP-LINK firmware, and 319 out of 330 in Xiaomi firmware.
As shown in the “Vulnerability/Patch Distinction” column, FoC-Sim has the potential to effectively distinguish vulnerable functions from patched functions. It correctly distinguishes 70 out of 97 functions in NETGEAR firmware, 142 out of 181 in TP-Link, and 289 out of 330 in Xiaomi. It demonstrates that FoC-Sim has the ability to overcome the challenge C3. However, we observe a failed case in Libgcrypt. With manual inspection, we find that there is a huge difference between the vulnerable functions from the firmware database and the vulnerability database in both text and structure. In other words, the vulnerable function in Libgcrypt has multiple significantly different versions, while our vulnerable database only includes one.
Compared to simple homology vulnerability detection, distinguishing between vulnerable functions and patched functions is a more difficult and highly specialized task, requiring more nuanced comparisons of binary code. Numerous studies have focused on binary code patch detection to tackle this challenge. For example, Fiber (zhang2018precise, ) generates binary-level patch signatures from source-level patches and performs a signature match between target and reference binaries using syntactic features such as control flow graphs (CFGs) and basic blocks. This process aligns instructions and generates symbolic constraints to identify patches. BinXray (xu2020patch, ) performs binary-level patch detection by extracting execution traces as signatures from binary functions, and matching them to the target function to confirm the presence of a patch. RoBin (yang2023towards, ) improves vulnerability verification through patch semantic analysis, providing a deeper understanding of how to exploit the semantic differences between vulnerable and patched versions to more accurately detect patches. As a binary code similarity calculation engine, FoC-Sim has the potential to achieve more effective patch detection capabilities in the future by incorporating the targeted innovative methods in patch detection from the above works.
Furthermore, we evaluate the ability of FoC-Sim to detect cryptographic functions in real-world stripped binaries. Specifically, we first use 1,336 binary files (containing 2,213,334 functions) from 14 projects collected in Table 2 as a Database Server of known cryptographic functions. Then, we select 92 unstripped binaries of 10 firmware from the constructed firmware database and annotate 2,000 cryptographic related function using the keyword-based discriminator mentioned in Section 3.3. Subsequently, we strip symbols from these binaries to generate fully stripped test samples to simulate real reverse analysis scenarios. Finally, we compare the similarity of these 2,000 functions with the function in the Database Server one by one and calculate their Recall@K metrics. Note that we ensure that each function has at least one homologous function in the Database Server. The experimental results, as shown in Table 14, indicate that when the retrieval range is small, FoC-Sim achieves a Recall@1 score of 0.714, meaning the model can accurately identify most cryptographic functions when faced with a smaller pool of candidates. As the retrieval range increases, FoC-Sim’s recognition ability significantly improves, particularly achieving a Recall@100 of 0.955, demonstrating its strong effectiveness in real-world cryptographic function detection scenarios. Removing different components of FoC-Sim leads to varying degrees of performance degradation, but it still maintains strong identification ability, with the code semantics being the most critical factor affecting detection performance.
@@
Answering RQ3: In the process of analyzing the encrypted communication module in the virus sample, FoC-BinLLM accurately predict the function names and summaries of related functions. FoC-Sim can effectively retrieve potential vulnerable cryptographic functions in the firmware, and correctly distinguish the subtle differences between the vulnerable functions and the patched versions. Furthermore, FoC-Sim can effectively detect cryptographic functions in real-world stripped binaries. In summary, the experimental results show the practical ability of FoC  in real-world scenarios.